OpenStudy (anonymous):

Multiply and simplify: x-6/x^2+4x-32 x-4/x+2

OpenStudy (anonymous):

what's with all the pretty pictures?

OpenStudy (anonymous):

i mean...questions :P

OpenStudy (anonymous):

can you put some brackets please? >_> lol

OpenStudy (pfenn1):

Is your problem \[\left[{(x-6)\over x^2+4x-32} \right] \times {(x-4) \over (x+2) }\]

OpenStudy (pfenn1):

or\[\left[ \left\{ (x-6) \over x^2 \right\} + 4x-32\right] \times \left[ (x-4) \over (x+2) \right]\]